OpenStudy (anonymous):

Solve the following system of equations:x + 2y + z = -2-2x + 4y - 4z = -222x + y + 2z = 5

OpenStudy (anonymous):

@jim_thompson5910 @Mertsj

OpenStudy (mertsj):

Could you please make a clearer post?

jimthompson5910 (jim_thompson5910):

is the actual problem this? x + 2y + z = -2 -2x + 4y - 4z = -22 2x + y + 2z = 5

OpenStudy (anonymous):

Solve the following system of equations:

OpenStudy (anonymous):

\[x + 2y + z = -2-2x + 4y - 4z = -222x + y + 2z = 5 \]

OpenStudy (anonymous):

That's how my worksheet has it written but it could be the way you have it @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's how the worksheet wrote it out? literally that format?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

well that's just lousy

jimthompson5910 (jim_thompson5910):

it's hard to tell where one equation ends and where one equation begins what I wrote out is just one possibility of what they meant

OpenStudy (anonymous):

I think so... it says under that the possible answers are:

OpenStudy (anonymous):

(3, -3, 1) (-3, -3, 1) (3, -3, -1) (-3, 3, 1)

jimthompson5910 (jim_thompson5910):

ok let me try out a few things real quick

jimthompson5910 (jim_thompson5910):

ok it has to the be that because the other possibilities give different answers

jimthompson5910 (jim_thompson5910):

Start off by solving the first equation for x x + 2y + z = -2 x + z = -2 - 2y x = -2 - 2y - z ------------------------------------------------------- Now plug x = -2 - 2y - z into the second equation -2x + 4y - 4z = -22 -2(-2 - 2y - z) + 4y - 4z = -22 4 + 4y + 2z + 4y - 4z = -22 4+8y-2z = -22 8y-2z = -22 - 4 8y-2z = -26 ------------------------------------------------------- Then plug x = -2 - 2y - z into the third equation 2x + y + 2z = 5 2( -2 - 2y - z) + y + 2z = 5 -4 - 4y - 2z + y + 2z = 5 -4-3y = 5

jimthompson5910 (jim_thompson5910):

From here, you need to solve -4-3y = 5 for y Once you have the value of y, you solve 8y-2z = -26 for z (after plugging in the value of y found previously)

jimthompson5910 (jim_thompson5910):

Finally, once you know y and z, you can use them to get x by plugging them into x = -2 - 2y - z